# IPPP/24/41

### v1

@@
August 2, 2024
Including exclusive heavy vector-meson
production data in a global parton analysis
@@
We perform a parton analysis within xFitter to determine the gluon parton distribution function (PDF) at next-to-leading order (NLO) at moderate-to-low $x$ using the measurements of exclusive $J/\psi$ production in $ep$ and $pp$ collisions from HERA and LHC. We further study the constraints from the corresponding $\Upsilon$ production process and finish by discussing the possible effects at next-to-next-to-leading order (NNLO) through incorporation of a $K$ factor for the exclusive heavy vector-meson coefficient function at NLO.
@@ 1 Introduction
The precision of the parton distributions of the proton is well established from global parton analyses provided that the values of the momentum fraction $x$ and resolution scales $Q^{2}$ are not too low. For $x>10^{-3}$, there is a strong agreement among the results from various global fit analyses [1, 2, 3]. However, as $x$ decreases, particularly at lower scales, the uncertainty in these distributions increases significantly. This increase in uncertainty is due to the lack of experimental data directly probing this region.
@@
In principle, this allows us to include the low-$x$ exclusive $J/\psi$ data in the global PDF analysis. In practice the problem is that the Shuvaev transform amounts to a slowly convergent double integral and the corresponding computation is time consuming. This is troublesome for the global fit as after every iteration, a new grid has to be computed to obtain the updated theory prediction.
@@ 2 Effective low-$x$ gluon PDF data
corresponding to energies $W_{\pm}$ depending on whether the vector meson travels in the direction of the photon $(W_{+})$, or against the direction of the photon $(W_{-})$. We see that the values of the subprocess cross sections $\sigma(\gamma p\to J/\psi p)$, for each rapidity $Y$ of the $J/\psi$, are weighted by the corresponding gap survival factors $S^{2}(W_{\pm})$ [11] and photon fluxes $dn/dk_{\pm}$ [12]. The cross section with the lower $\gamma p$-energy ($W_{-}$) was measured at HERA [13, 14, 15, 16] and can be calculated with sufficiently good accuracy since it corresponds to relatively large $x$ where the uncertainties in the parton distributions are small. Hence the exclusive $J/\psi$ data provide reliable values for $\sigma_{+}(\gamma+p\to J/\psi+p)$, corresponding to the component with the larger $\gamma p$-energy, $W_{+}$, see Fig. 3.
@@ 7 TeV
Table 1: The values of $xg_{\rm eff}$ and $\delta xg_{\rm eff}$ obtained from eqns. (6) and (7) using the exclusive $J/\psi$ production data at 7 TeV [4] and 13 TeV [5] obtained by the LHCb collaboration.
@@
Table 2: The values of $xg_{\rm eff}$ and $\delta xg_{\rm eff}$ obtained from eqns. (6) and (7) using the exclusive $\Upsilon$ production data at 7, 8 TeV [6] obtained by the LHCb collaboration.
In Table 2, we show the final effective iterated gluon points for exclusive $\Upsilon$ production [6] at the corresponding ‘optimal’ scale 22.4 GeV2. The accuracy of these points is worse than that obtained from $J/\psi$ production, but they allow for a constraint on the $Q^{2}$-evolution in the low-$x$ region. [^6] So far the exclusive $\Upsilon$ data for the $\gamma p\to\Upsilon p$ cross section from the LHCb collaboration have been obtained using gap survival factors $S^{2}(W_{\pm})$ and photon fluxes $dn/dk_{\pm}$ that need updating or correcting, see [22]. Nevertheless, the resulting $\Upsilon$ data, shown in Table 2 already include the corresponding corrections, which however never exceed the experimental error bars. Note that the resulting $\Upsilon$ points already show evidence of a rising gluon PDF power behaviour in the region $10^{-5}<x<10^{-4}$ and $Q^{2}=m_{b}^{2}\sim 22~{}{\rm GeV}^{2}$ (see Fig. 2 of [22]).
@@ 3 The impact of the low-$x$ gluon effective data points on a global analysis
It is clearly seen in Fig. 4 that the inclusion of the low-$x$ effective gluon points extracted from the exclusive heavy vector-meson production at the LHC essentially improve the accuracy of the global parton analysis, and we obtain a good overall final $\chi_{\text{min}}^{2}/\text{d.o.f}\sim O(1)$. The blue error band is much smaller than the red one. [^7] As usual, additional theoretical uncertainties such as the correction to the NRQCD approach used for the $q\bar{q}\rightarrow V$ transition in the computation of $\sigma(\gamma+p\to V+p)$, possible higher twist, effects of absorptive corrections or other higher-order corrections are not shown. Recall that taking in its present form our approach satisfactorily describes the available HERA data [17]. Moreover, the inclusion of new information obtained from the heavy vector-meson ultraperipheral production affects (albeit expectedly not so strongly) even the quark distribution as demonstrated, for example, in Fig. 5.
@@ 5 Conclusion and outlook
Since at present we do not have the NNLO coefficient functions for the photon to vector meson ($\gamma\to V$) transition, our computations were performed at NLO level. However, there is the possibility to reach NNLO* accuracy, by extracting the NNLO/NLO $K$ factor from the existing HERA $\gamma+p\to J/\psi+p$ data as discussed in Section 4. Moreover, exclusive vector-meson production data have already been collected by the LHC and more should follow. Thus further theoretical study of the NNLO* analysis is merited.
This current analysis has already demonstrated the potential of the existing data to significantly improve the NLO global parton distributions at small $x$.

### v2

@@
January 28, 2025
A method to include exclusive heavy vector-meson
production data at small $x$ in global parton analyses
@@
We propose a method which allows the inclusion of exclusive heavy vector-meson production data at low $x$ in future global parton analyses. As an example we perform a study within xFitter to determine the gluon parton distribution function (PDF) at next-to-leading order (NLO) at moderate-to-low $x$ using the measurements of exclusive $J/\psi$ production in $ep$ and $pp$ collisions from HERA and LHC. We further study the constraints from the corresponding $\Upsilon$ production process. We finish by discussing the possible effects at next-to-next-to-leading order (NNLO) through incorporation of a $K$ factor for the exclusive heavy vector-meson coefficient function at NLO.
@@ 1 Introduction
[…]
@@
In principle, this allows us to include the low-$x$ exclusive $J/\psi$ data in the global PDF analysis. In practice the problem is that the Shuvaev transform amounts to a slowly convergent double integral and the corresponding computation is time consuming. This is troublesome for the global fit as after every iteration, a new three-dimensional grid defined over the variables $X,\xi$ and $Q^{2}$ has to be computed to obtain the updated theory prediction.
In the present paper we describe a method which helps to overcome this problem by replacing the exclusive vector-meson data at small $x$ by ‘effective gluon points’. To demonstrate the efficiency of the proposed method and to show how the inclusion of the LHCb $J/\psi$ and $\Upsilon$ exclusive cross sections affect the result we compare the standard xFitter [11] NLO parton analysis of DIS data with that supplemented by the vector-meson data. Note that our goal is not to present a new set of parton distributions but to emphasise the possible role of the vector-meson data in global parton analyses. The new ‘effective gluon points’ presented in Tables 1 and 2 can be used in future global analyses.
@@ 2 Effective low-$x$ gluon PDF data
corresponding to energies $W_{\pm}$, where $W_{+}$ denotes the energy when the vector meson travels in the same direction as the photon, and $W_{-}$ when it travels in the opposite direction. We see that the values of the subprocess cross sections $\sigma(\gamma p\to J/\psi p)$, for each rapidity $Y$ of the $J/\psi$, are weighted by the corresponding gap survival factors $S^{2}(W_{\pm})$ [12] and photon fluxes $dn/dk_{\pm}$ [13]. The cross section with the lower $\gamma p$-energy ($W_{-}$) was measured at HERA [14, 15, 16, 17] and can be calculated with sufficient accuracy since it corresponds to relatively large $x$ where the uncertainties in the parton distributions are small. Hence the exclusive $J/\psi$ data provide reliable values for $\sigma_{+}(\gamma+p\to J/\psi+p)$, corresponding to the component with the larger $\gamma p$-energy, $W_{+}$, see Fig. 3.222Actually, the subtraction of the $W_{-}$ contribution was done already in the LHCb analyses where the results are presented as the $\sigma_{+}(\gamma+p\rightarrow V+p)$ cross sections.
@@ 7 TeV
Table 1: The values of $xg_{\rm eff}$ and $\delta xg_{\rm eff}$ obtained from eqns. (6) and (7) at $Q^{2}=M_{J/\psi}^{2}/4=2.4\,{\rm GeV}^{2}$ using the exclusive $J/\psi$ production data at 7 TeV [4] and 13 TeV [5] obtained by the LHCb collaboration.
@@
Table 2: The values of $xg_{\rm eff}$ and $\delta xg_{\rm eff}$ obtained from eqns. (6) and (7) at $Q^{2}=M_{\Upsilon}^{2}/4=22.4\,{\rm GeV}^{2}$ using the exclusive $\Upsilon$ production data at 7, 8 TeV [6] obtained by the LHCb collaboration.
In Table 2, we show the final effective iterated gluon points for exclusive $\Upsilon$ production [6] at the corresponding ‘optimal’ scale 22.4 GeV2. The accuracy of these points is worse than that obtained from $J/\psi$ production, but they allow for a constraint on the $Q^{2}$-evolution in the low-$x$ region. [^7] So far the exclusive $\Upsilon$ data for the $\gamma p\to\Upsilon p$ cross section from the LHCb collaboration have been obtained using gap survival factors $S^{2}(W_{\pm})$ and photon fluxes $dn/dk_{\pm}$ that need updating or correcting, see [24]. Nevertheless, the resulting $\Upsilon$ data shown in Table 2 already include the corresponding corrections, which however never exceed the experimental error bars. Note that the resulting $\Upsilon$ points already show evidence of a rising gluon PDF power behaviour in the region $10^{-5}<x<10^{-4}$ and $Q^{2}=m_{b}^{2}\sim 22~{}{\rm GeV}^{2}$ (see Fig. 2 of [24]). Recall that as was shown in [24], at the border of the LHCb acceptance, the choice of survival factor and photon flux only amounted to 5% for $J/\psi$ production, while for $\Upsilon$ production it is of the order of 25%.
For the results presented in Tables 1 & 2, we actually use the Shuvaev transform for $x\leq 10^{-4}$ ($x\leq 10^{-3}$ for the data presented in Fig. 3).
@@ 3 The impact of the low-$x$ gluon effective data points on a global analysis
It is clearly seen in Fig. 4 that the inclusion of the low-$x$ effective gluon points extracted from the exclusive heavy vector-meson production at the LHC essentially improve the accuracy of the global parton analysis, and we obtain a good overall final $\chi_{\text{min}}^{2}/\text{d.o.f}\sim O(1)$. The blue error band is much smaller than the red one. [^8] As usual, additional theoretical uncertainties such as the correction to the Non-Relativistic QCD (NRQCD) approach used for the $q\bar{q}\to V$ transition in the computation of $\sigma(\gamma p\to V+p)$, or possible higher twist effects, or effects of absorptive corrections, or other higher-order corrections, are not shown. Note that, apart from the absorptive effects, all other corrections (e.g. NRQCD) do not affect the energy behaviour of the amplitude, but just change its normalisation. Indeed, to ‘know’ the initial energy we need an additional gluon, which connects the upper (photon block in Fig. 1) with the lower/target parts of the diagram. This would correspond to a higher-twist contribution and should be suppressed by a large factorisation scale $\mu_{F}$. Typically, these corrections act ‘locally’ in rapidity space and therefore do not affect the energy behaviour of the amplitude, only its normalisation. On the other hand, our approach, in its present form, satisfactorily describes the available HERA data [18]. Recall that, as was shown in [19], normalising the vector-meson wave function to the experimentally measured $J/\psi\to l^{+}l^{-}$ decay width allows us to account for relativistic corrections in the NRQCD approach with good accuracy (within a few percent). Moreover, the inclusion of new information obtained from the heavy vector-meson ultraperipheral production affects (albeit expectedly not so strongly) even the quark distribution as demonstrated, for example, in Fig. 5. The new $g_{\text{eff}}$ low-$x$ points affect the quark distribution even at a rather large $x$ due to the energy-momentum sum rules and the particular choice of the fixed form of the input ansatz.
@@ 5 Conclusion and outlook
Since at present the NNLO coefficient functions for the photon to vector meson ($\gamma\to V$) transition are not available, our computations were performed at NLO level. However, there is the possibility to reach NNLO* accuracy, by extracting the NNLO/NLO $K$ factor from the existing HERA $\gamma+p\to J/\psi+p$ data as discussed in Section 4. Moreover, exclusive vector-meson production data have already been collected by the LHC and more should follow. This merits further theoretical study of the NNLO* analysis, which we reserve for future work.
The analysis presented in this work has demonstrated the potential of existing exclusive data to significantly improve the NLO global parton distributions at small $x$.